Two tennis balls are dropped through the air from a tall building. One of them is filled with lead  pellets making it heavier than the other. In presence of air-resistance, the ball to reach the ground first is the ----

a. Lead filled tennis ball

b. Insufficient infomation

c. Regular tennis ball

d. both hit the ground at the same time

2. As you drive in your car at 22.5 mi/h, you see a child’s ball roll into the street ahead of you. You hit the brakes and stop as quickly as you can. In this case, you come to rest in 2 s. How far does your car travel as you brake to a stop? (Use: 1 mile= 1600 m, 1 h= 3600s).

a. About 22 m

b. About 10 m

c. About 10 miles

d. About 1.2 m
Answer #1

1)
a)

Lead filled tennis ball will reach the ground first.

The lead filled tennis ball and the normal tennis ball will both experience the same air resistance, but as the lead filled ball is heavier the acceleration of the lead filled ball will be higher.

acceleration_{lead\ filled}=\frac{Mg-(Air\ resistance)}{M}=g-\frac{(Air\ resistance)}{M}

acceleration_{normal}=\frac{mg-(Air\ resistance)}{m}=g-\frac{(Air\ resistance)}{m}

Since M>m,

acceleration_{lead\ filled}>acceleration_{normal}

2)

v=u+at

0=22.5*\frac{1600}{3600}+a*2

a=-5\ m/s^{2}

2as=v^{2}-u^{2}

2*(-5)*s=0^{2}-(22.5*\frac{1600}{3600})^{2}

s=10\ m

option b) is the right answer